3. Suppose that X and Y are related by the simple linear regression model Y = a + BX +E where a, 8 are unknown parameters, an(b) (10 points) Compute the first and the second moment of B (in terms of a,B,0).(c) (10 points) Compute the first and the second moment of â in terms of a, ß,0).

Help
Answer #1

A ML C Sel7, . (4) - * , * * , + + 1 + Ks - 1 + 2 + 3 + 1 + 5 2 เร เ (๔ +- เ 8, ) เ 5 , 6 : : : () (21 เ | เ: 1 0 0 I mean of5 x 7. L :( * , ( x + B + 8. -| 2 a{nit B E xixi +Exiei = 42 ส. - 22 12 +2 8. - เรณ + 5S B + 2 + 222 + 362 ฯ S9+ 555Ŕ by first moment of Å is the expected value of which is given by the Expression E ( 8 = 2 ( x. - ) = (Y) {? - ne? = {chi-noco is nothing but the second mom variance of ê = Ža: -2) Yi ( **-) var (ês = Z (hi-x² Var (Y;) (x2- não)? - 02 E (Mina) 2 (%.- co first moment of expected value â is the â= ý - ² â = { g; - sê Ela) = E(Y) - ã ECB) n - Ela+ pai) - ã B n = - a + ßñ - aa second moment of ê is the variance of å ý ê var câr= 02e ne? nExi - ne?) o2x55 (1 5 ( 55 - 5*32) 1152 ans 11 ? x 1l 10 Note